# EVA-Gaussian: 3D Gaussian-based Real-time Human Novel View Synthesis under Diverse Camera Settings

### v1

@@ Abstract
The feed-forward based 3D Gaussian Splatting method has demonstrated exceptional capability in real-time human novel view synthesis. However, existing approaches are restricted to dense viewpoint settings, which limits their flexibility in free-viewpoint rendering across a wide range of camera view angle discrepancies. To address this limitation, we propose a real-time pipeline named EVA-Gaussian for 3D human novel view synthesis across diverse camera settings. Specifically, we first introduce an Efficient cross-View Attention (EVA) module to accurately estimate the position of each 3D Gaussian from the source images. Then, we integrate the source images with the estimated Gaussian position map to predict the attributes and feature embeddings of the 3D Gaussians. Moreover, we employ a recurrent feature refiner to correct artifacts caused by geometric errors in position estimation and enhance visual fidelity. To further improve synthesis quality, we incorporate a powerful anchor loss function for both 3D Gaussian attributes and human face landmarks. Experimental results on the THuman2.0 and THumansit datasets showcase the superiority of our EVA-Gaussian approach in rendering quality across diverse camera settings. Project page: https://zhenliuzju.github.io/huyingdong/EVA-Gaussian.
@@ 1 Introduction
3D reconstruction and novel view synthesis have long been fundamental yet complex tasks in visual data representation and computer vision. Recent advancements in fast 3D reconstruction and novel view synthesis for humans have shown immense potential in applications such as holographic communication, real-time teaching, and augmented/virtual reality (AR/VR), where time efficiency is critical for user experience and downstream processing. Nonetheless, existing methods either rely on dense input views and precise templates as prior knowledge (Qian et al., 2024; Lei et al., 2024; Hu et al., 2024; Wen et al., 2024; Kocabas et al., 2024; Kwon et al., 2024) or are restricted to specific camera poses (Zheng et al., 2024; Tu et al., 2024). None of these approaches have fully developed a pipeline for real-time human reconstruction under diverse, especially sparse, camera viewpoints, which remains a significant challenge.
@@
In recent years, Neural Radiance Fields (NeRFs) (Mildenhall et al., 2021) have emerged as a promising technique for 3D reconstruction. These models employ neural networks to predict the color and density of sampled 3D points along camera rays and aggregate these predictions to synthesize novel images with high fidelity. Despite their effectiveness, NeRFs suffer from substantial time consumption during both the training and rendering phases. Although various advancements, such as multi-resolution hash encoding (Müller et al., 2022) and feed-forward neural scene prediction (Yu et al., 2021a; Xu et al., 2024), have been made to reduce the time for reconstruction and rendering, the achievable speeds remain insufficient for real-time applications.
@@
More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has been introduced as a solution to this rendering bottleneck. 3DGS utilizes a set of discrete Gaussian representations to model complex 3D scenes and leverages the $\alpha$-blending technique to enable real-time novel view rendering. In the field of 3D human avatar reconstruction, previous works (Qian et al., 2024; Lei et al., 2024; Hu et al., 2024; Wen et al., 2024; Kocabas et al., 2024; Kwon et al., 2024) have employed 3DGS as a representation for humans to achieve animatable full-body human avatar reconstruction. These methods, however, rely on precise human templates as priors, and their training and reconstruction processes can take from minutes to hours, which impedes their use in real-time applications such as holographic communication. While a feed-forward human reconstruction method (Zheng et al., 2024) has achieved fast reconstruction and real-time rendering with stereo inputs, the stereo settings and small angle change between camera viewpoints restrict the overall reconstruction quality under sparse camera settings and lead to sub-optimal performance with more than two input views.
@@
To address these limitations and enable real-time 3D reconstruction of humans using various camera positions and different numbers of cameras, we propose EVA-Gaussian, a novel 3D Gaussian-based pipeline for real-time human novel view synthesis. Our method attaches 3D Gaussians to the surface of human body through multi-view depth estimation and aligns their positions closely with point cloud locations. A key innovation of our method is the introduction of an Efficient cross-View Attention (EVA) module for multi-view 3D Gaussian position estimation (see Sec. 4.2). Specifically, we employ a U-Net (Ronneberger et al., 2015) as the backbone and further use dedicated window-embedded cross-view attention to infer multi-view position correspondences. This attention mechanism enables EVA-Gaussian to effectively process multiple inputs from cameras positioned at various viewpoint angles, thereby ensuring robust performance across a wide range of viewing angles, even under extremely sparse camera settings. Besides, we incorporate a Gaussian attribute estimation module that takes the EVA output and the original RGB images as input to estimate the remaining 3D Gaussian attributes (see Sec. 4.3). Furthermore, we embed an additional attribute, referred to as feature, into each Gaussian for further feature splatting and image quality refinement, thereby mitigating the position estimation error introduced by the EVA module (see Sec. 4.4). In addition, we employ an anchor loss to penalize the inconsistency between multi-view face landmarks, which achieves better supervision for human faces (see Sec. 4.5). We conduct extensive experiments on the THuman2.0 (Yu et al., 2021b) and THumanSit (Zhang et al., 2023) datasets. The results, as exemplified in Fig. 1, demonstrate that our proposed EVA-Gaussian outperforms existing feed-forward synthesis approaches in rendering quality, while enabling real-time reconstruction and rendering. Moreover, our approach generalizes well to settings with diverse numbers of cameras and significant changes in camera viewpoint angles. In summary, our main contributions are as follows:
@@
  • We propose a novel pipeline for fast feed-forward 3D human reconstruction, called EVA-Gaussian, that comprises three main stages: 1) a multi-view 3D Gaussian position estimation stage, 2) a 3D Gaussian attributes estimation stage, and 3) a feature refinement stage.
@@
  • We employ a recurrent feature refiner that fuses splatted RGB images and feature maps to mitigate geometric artifacts caused by position estimation errors. Moreover, we incorporate an anchor loss that utilizes facial landmarks as anchor points to better supervise Gaussian position estimation, thereby enhancing the quality of synthesized novel view images.
  • Extensive experiments on THuman2.0 and THumansit demonstrate the effectiveness and superiority of our proposed pipeline over existing methods in terms of rendered novel view quality and inference speed, especially under sparse camera settings.
@@ 2 Related Works
3DGS-based Human Reconstruction. 3D Gaussians Splatting has recently emerged as an effective technique for 3D human reconstruction. However, most previous works (Lei et al., 2024; Hu et al., 2024; Wen et al., 2024; Qian et al., 2024; Kocabas et al., 2024) bind 3D Gaussians to a predefined human mesh model, such as SMPL (Loper et al., 2023) or SMPL-X (Pavlakos et al., 2019). This approach generates 3D Gaussians and human models in a canonical space and then transforms them to match the target human pose using the predefined weights of the human model. This iterative binding process, however, is extremely time-consuming. Moreover, these methods require human templates as inputs at each frame, which incurs extra computational cost and potentially misleads the reconstruction procedure due to the errors in pose estimation. These limitations significantly hinder their applicability in real-world scenarios.
@@
Fast Generalizable 3D Reconstruction. In the field of NeRF rendering, pixelNeRF (Yu et al., 2021a) pioneers the approach of predicting features per pixel from a single image in a feed-forward manner for 3D reconstruction. While subsequent works (Chen et al., 2021; Wang et al., 2021; Lin et al., 2022) have followed this feed-forward NeRF pipeline, they still suffer from the extensive time consumption of the NeRF rendering process. Besides, their reconstruction results are often unsatisfactory in sparse camera settings. The introduction of 3DGS has helped mitigate the rendering speed issue of high-quality novel view synthesis methods. Notably, pixelSplat (Charatan et al., 2024) and Splatter Image (Szymanowicz et al., 2024) are the first to combine the feed-forward inference and 3DGS, which predict 3D Gaussian attributes for each pixel and project them back to the 3D space for novel view synthesis in a real-time manner. Nevertheless, these methods still struggle with inaccurate estimation of 3D Gaussian positions. MVSplat (Chen et al., 2024) and MVGaussian (Liu et al., 2024a) address this issue by leveraging cost-volume modules, thereby achieving better novel view image quality. Moreover, latentSplat (Wewer et al., 2024) attaches a latent vector to each 3D Gaussian and uses this vector for novel view refinement through a diffusion decoder and generative loss, which significantly improves image quality on extrapolation views. Despite these advancements, these approaches do not fully exploit prior knowledge about human images and camera settings, which limits their performance on human reconstruction and novel view synthesis. In the field of 3D generation, LGM (Tang et al., 2024) and GSLRM (Zhang et al., 2024) employ a transformer-based feed-forward pipeline for predicting 3D Gaussian attributes from multi-view images in real-world scenarios, but these networks are often too complex for real-time reconstruction. Our work utilizes a carefully designed memory-efficient attention mechanism, thereby enabling generalizable 3D reconstruction at high resolutions with a much lower temporal and computational cost.
@@
The work most closely related to ours is GPS-Gaussian (Zheng et al., 2024), which proposes a stereo matching network for 3D Gaussian position estimation and employs two 3-layer U-Nets to predict 3D Gaussian scales, rotations, and opacities. Although GPS-Gaussian has demonstrated the capability for real-time human reconstruction and novel view synthesis, it suffers from severe distortions under sparse camera settings and mismatch across multiple viewpoints. Subsequent works have attempted to alleviate these issues. For instance, Tele-Aloha (Tu et al., 2024) introduces an image blending and cascaded disparity estimation method for human reconstruction with four input views. However, this approach is tailored to a specific system and struggles to generalize to sparser camera settings. On the other hand, GHG (Kwon et al., 2024) achieves real-time 3D Gaussian-based human novel view synthesis in a feed-forward manner, but it requires additional human template priors, thus inheriting the limitations associated with template-based human reconstruction methods. In contrast, our proposed method eliminates the need for human templates and is specifically designed to generalize effectively across various sparse camera settings.
3 Preliminary
3D Gaussian Splatting (3DGS) uses a set of 3D Gaussian distributions to represent a 3D sence based on a sequence of multi-view RGB images (Kerbl et al., 2023). Each 3D Gaussian can be mathematically expressed as:
$G(\bm{x})=e^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})},$ (1)
where $\bm{\mu}\in\mathbb{R}^{3}$ and $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$ denote the position vector and the covariance matrix, respectively. The covariance matrix $\bm{\Sigma}$ is constrained to be semi-positive definite, which is guaranteed by decomposing it into a rotation matrix $\bm{R}\in\mathbb{R}^{3\times 3}$ and a scaling matrix $\bm{S}\in\mathbb{R}^{3\times 3}$ as:
$\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{T}\bm{R}^{T},$ (2)
where $\bm{R}$ is further represented by a quaternion $\bm{q}\in\mathbb{R}^{4}$, and $\bm{S}$ is further represented by a scaling vector $\bm{s}\in\mathbb{R}^{3}$. In addition, to facilitate rendering from a specific camera perspective, 3DGS leverages a view transformation $\bm{W}$ to transfer the 3D Gaussians from the world space to the camera space, and employs a projective transformation $\bm{J}$ to approximate the projection of these 3D Gaussians onto the 2D image plane.
The final rendered color for each pixel is computed using an $\alpha$-blending function that accumulates the contributions of all the projected 3D Gaussians on the pixel:
$\bm{c}_{\text{pixel}}=\sum_{i=1}^{N}\bm{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
where $N$ denotes the number of Gaussians, $\alpha_{i}$ is the product of the $i$-th projected 2D Gaussian and its optimized opacity value $o_{i}\in[0,1]$, and $\bm{c}_{i}\in\mathbb{R}^{3}$ denotes the color attribute of the Gaussian, which is represented using spherical harmonics coefficients to enable view-dependent color modeling.
@@ 4.1 Overview
where $\bm{\theta}$ denotes the parameters of the learnable neural network.
The framework of EVA-Gaussian is depicted in Fig. 2. EVA-Gaussian splits the process of predicting Gaussian maps into three stages. In the first stage, it employs a U-Net architecture with an Efficient cross-View Attention module (EVA) to obtain enhanced multi-view predictions of the 3D Gaussian position maps $\{\bm{P}_{i}\}_{i=1}^{n}$, as elaborated in Sec. 4.2. In the second stage, a Gaussian attribute prediction network, detailed in Sec. 4.3, takes the predicted 3D Gaussian position maps $\{\bm{P}_{i}\}_{i=1}^{n}$ and the original RGB images $\{\bm{I}_{i}\}_{i=1}^{n}$ as input to estimate the remaining attributes of 3D Gaussians. The predicted 3D Gaussians from all source images are then aggregated to render target views using differential rasterization (Kerbl et al., 2023). In the final stage, the rendered RGB image $\hat{\bm{I}}^{0}$ and its corresponding feature map $\bm{F}_{\text{novel}}$ are fused for further refinement using the network described in Sec. 4.4. In addition, an anchor loss is introduced during the training stage to enhance the overall reconstruction quality, as depicted in Sec. 4.5.
@@ 4.2 Gaussian Position Estimation
To tackle this challenge, we employ a U-Net based architecture $\mathcal{D}^{P}_{\bm{\theta}_{1}}$ to estimate the 3D Gaussian position maps $\{\bm{P}_{i}\}_{i=1}^{n}$ from multi-view images $\{\bm{I}_{i}\}_{i=1}^{n}$, which is expressed as:
@@
To ensure accurate depth estimation across diverse camera angles or arbitrary input views, we propose an EVA module, as illustrated in Fig. 3. This module is integrated into the three lowest resolution layers of the U-Net basebone $\mathcal{D}^{P}_{\bm{\theta}_{1}}$ to facilitate the multi-view correspondence retrieval and information exchange. We use $j$ to denote the index of each of these three layers, with $j=-1,j=-2,$ and $j=-3$ representing the lowest, the second-lowest, and the third-lowest resolution layers, respectively. EVA takes multiple intermediate image features $\bm{E}^{j}_{i}\in\mathbb{R}^{R^{j}\times C^{j}},\forall i\in\{1,\cdots,n\},\forall j\in\{-1,-2,-3\}$, as input and outputs the corresponding enhanced image features $\widetilde{\bm{E}}^{j}_{i}$, where $R^{j}$ and $C^{j}$ denote the total number of pixels and the channel dimension of each pixel at layer $j$, respectively. Before the execution of attention mechanisms, a learnable positional embedding $\gamma$ is added to the intermediate feature $\bm{E}^{j}_{i}$ to improve the understanding of image coordinates.
Figure 3: Efficient cross-View Attention (EVA) module for 3D Gaussian position estimation. EVA takes multi-view image features as input, embeds them into window patches using a shifted algorithm, and performs cross-view attention between the features from different views.
In contrast to other feed-forward scene reconstruction methods (Chen et al., 2024; Charatan et al., 2024; Liu et al., 2024a) that apply to a low resolution of 256$\times$256, our approach aims for high-quality human reconstruction at 1024 resolution. Given that the corresponding pixels from the reference views are located in adjacent locations only under human-centric camera settings, calculating attention scores across a concatenated multi-view image sequence or feature map for the entire image, as adopted by Chen et al. (2024), is highly inefficient. To improve efficiency, EVA computes cross-attention only within a local window, which is shifted by half the window size at each attention iteration. This shifted window embedding allows for low computation complexity and better GPU memory utilization, while maintaining high performance.
[…]
where $\bm{A}^{j}_{i},\forall i\in\{1,2,\cdots,n\}$, denotes the resultant attention output.
Notably, when the scale of each Gaussian is sufficiently small, the 3D Gaussian position of a pixel aligns precisely with its corresponding value on the depth map. A detailed proof of this property is provided in Appendix C. Based on this observation, we train the position estimation network $\mathcal{D}^{P}_{\bm{\theta}}$ to obtain the position maps $\{\bm{P}_{i}\}_{i=1}^{n}$ with the mean squared error (MSE) loss function:

### v2

# EVA-Gaussian: 3D Gaussian-based Real-time Human Novel View Synthesis under Diverse Multi-view Camera Settings

@@ Abstract
Feed-forward based 3D Gaussian Splatting methods have demonstrated exceptional capability in real-time novel view synthesis for human models. However, current approaches are confined to either dense viewpoint configurations or restricted image resolutions. These limitations hinder their flexibility in free-viewpoint rendering across a wide range of camera view angle discrepancies, and also restrict their ability to recover fine-grained human details in real time using commonly available GPUs. To address these challenges, we propose a novel pipeline named EVA-Gaussian for 3D human novel view synthesis across diverse multi-view camera settings. Specifically, we first design an Efficient Cross-View Attention (EVA) module to effectively fuse cross-view information under high resolution inputs and sparse view settings, while minimizing temporal and computational overhead. Additionally, we introduce a feature refinement mechianism to predict the attributes of the 3D Gaussians and assign a feature value to each Gaussian, enabling the correction of artifacts caused by geometric inaccuracies in position estimation and enhancing overall visual fidelity. Experimental results on the THuman2.0 and THumansit datasets showcase the superiority of EVA-Gaussian in rendering quality across diverse camera settings.
@@ 1 Introduction
3D reconstruction and novel view synthesis have long been fundamental yet complex tasks in visual data representation and computer vision. Recent advancements in fast 3D reconstruction and novel view synthesis for humans have shown immense potential in applications such as holographic communication, real-time teaching, and augmented/virtual reality (AR/VR), where time efficiency and the clarity of human body representations are critical for delivering satisfactory user experiences. Nonetheless, existing methods either rely on dense input views and precise templates as prior knowledge [21, 9, 5, 29, 7, 8] or are restricted to specific camera poses [36, 26]. None of these approaches has fully developed a pipeline for real-time human reconstruction under diverse, especially sparse, camera viewpoints while maintaining high-resolution input quality.
@@
In recent years, Neural Radiance Fields (NeRFs) [17] have emerged as a promising technique for 3D reconstruction. These models employ neural networks to predict the color and density of sampled 3D points along camera rays and aggregate these predictions to synthesize novel images with high fidelity. Despite their effectiveness, NeRFs suffer from substantial time consumption during both the training and rendering phases. Although various advancements, such as multi-resolution hash encoding [18] and feed-forward neural scene prediction [32, 31], have been achieved to mitigate these inefficiencies, the achievable speeds remain insufficient for real-time applications.
@@
More recently, 3D Gaussian Splatting (3DGS) [6] has been introduced as a solution to the rendering bottleneck. 3DGS utilizes a set of discrete Gaussian representations to model complex 3D scenes and leverages the $\alpha$-blending technique to enable real-time novel view synthesis. In the field of 3D human avatar reconstruction, previous works [21, 9, 5, 29, 7, 8] have employed 3DGS as a representation for humans to achieve animatable full-body human avatar reconstruction. These methods, however, rely on precise human templates as priors, and their training and reconstruction processes can span from minutes to hours, which impedes their use in real-time applications such as holographic communication. While a feed-forward human reconstruction method [36] has achieved fast reconstruction and real-time rendering with stereo inputs at a resolution of 1024, its reliance on stereo settings and limited camera angle variations restricts its reconstruction quality under sparse camera settings and lead to sub-optimal performance when more than two input views are utilized.
@@
To address these limitations and enable high-resolution real-time 3D reconstruction of humans across diverse camera positions and varying numbers of cameras, we propose EVA-Gaussian, a novel 3D Gaussian-based pipeline for real-time human novel view synthesis. Our method attaches 3D Gaussians to the surface of the human body through multi-view depth estimation and aligns their positions closely with point cloud locations. A key innovation is the introduction of an Efficient cross-View Attention (EVA) module for multi-view 3D Gaussian position estimation (see Sec. 3.2). Specifically, we employ a U-Net [22] as the backbone and incorperate a dedicated window-embedded cross-view attention mechanism to infer multi-view position correspondences. This design leverages the observation that, in most human capture systems, human subjects are typically centered within each frame. As a result, the epipolar lines tend to align nearly parallel to the x-axis of the image, which enables us to significantly reduce the memory for multi-view attention while maintaining accuracy under large view discrepancies. Besides, we incorporate a Gaussian attribute estimation module that takes the EVA output and the original RGB images as input to estimate the remaining 3D Gaussian attributes (see Sec. 3.3). Furthermore, we embed an additional attribute, referred to as feature, into each Gaussian for further feature splatting and image quality refinement, thereby mitigating the position estimation errors introduced by the EVA module (see Sec. 3.4). In addition, we introduce an anchor loss to regularize the opacity and scale attributes of the 3D Gaussians (see Sec. 3.5). This ensures consistency between the point cloud depth and the 3D Gaussian position map, enhancing overall stability and accuracy. We conduct extensive experiments on the THuman2.0 [33] and THumanSit [34] datasets. The results, as exemplified in Fig. 1, demonstrate that EVA-Gaussian outperforms existing feed-forward synthesis approaches in rendering quality, while enabling real-time reconstruction and rendering. Moreover, our approach generalizes well to settings with varying numbers of cameras and significant changes in camera viewpoint angles. Our main contributions are summarized as follows:
@@
  • We propose EVA-Gaussian, a novel pipeline for fast feed-forward 3D human reconstruction that comprises three main stages: 1) 3D Gaussian position estimation, 2) 3D Gaussian attributes estimation, and 3) feature refinement.
@@
  • We embed a feature value to each 3D Gaussian through a feed-forward neural network and employ a recurrent feature refiner that fuses splatted images and feature maps to mitigate artifacts caused by position estimation errors.
  • Extensive experiments on the THuman2.0 and THumansit datasets demonstrate the effectiveness and superiority of our proposed pipeline over existing methods in terms of rendered novel view quality and inference speed, especially under sparse camera settings.
@@ 2 Related Works
Instant Human Reconstruction. Recovering a 3D human model from sparse view inputs and generating novel views in real time has been a long-standing challenge in computer vision and graphics. Despite continuous advancements, achieving both accuracy and efficiency remains an unresolved problem. PIFu [23] is one of the first methods to successfully reconstruct the human surface and its color map from one or several RGB images, leveraging the strong spatial encoding capabilities of implicit function representations. However, it remains constrained by low speed and resolution. Subsequent works attempt to address these limitations by employing specific rendering methods [10, 24] or incorperating additional depth information [33]. Nonetheless, their reliance on traditional 3D representations results in suboptimal rendering quality and efficiency.
@@
3DGS-based Human Avatar Reconstruction. 3D Gaussian Splatting has recently emerged as an effective technique for 3D human reconstruction. However, most previous works [9, 5, 29, 21, 7, 19] bind 3D Gaussians to a predefined human mesh model, such as SMPL [14] or SMPL-X [20]. This approach generates 3D Gaussians and human models in a canonical space and then transforms them to match the target human pose using predefined weights. This iterative binding process, however, is extremely time-consuming. Moreover, these methods require human templates as inputs at each frame, which incurs extra computational costs and potentially misleads the reconstruction due to errors in pose estimation. These limitations significantly hinder their applicability in real-world scenarios.
@@
Fast Generalizable 3D Reconstruction. In the field of NeRF rendering, pixelNeRF [32] pioneers the approach of predicting per-pixel features from a single image in a feed-forward manner for 3D reconstruction. While subsequent works [3, 27, 11] have followed this feed-forward NeRF pipeline, they still suffer from the extensive time consumption of the NeRF rendering process. Besides, their reconstruction results are often unsatisfactory under sparse camera settings. The introduction of 3DGS has helped mitigate the rendering speed issue of high-quality novel view synthesis. Notably, pixelSplat [2] and Splatter Image [25] are the first to combine feed-forward inference with 3DGS, which predict 3D Gaussian attributes for each pixel and project them back into the 3D space for real-time novel view synthesis. Nevertheless, they still struggle with inaccurate estimation of Gaussian positions. MVSplat [4] and MVGaussian [12] address this issue by leveraging cross-attention mechanisms and cost-volume modules, thereby achieving superior novel view quality. Moreover, latentSplat [30] attaches a latent vector to each 3D Gaussian and refines novel views through a diffusion decoder and generative loss, significantly improving image quality in extrapolation views. Despite these advancements, existing methods fail to fully exploit prior knowledge about human images and camera settings, which limits their performance on real-time human reconstruction and novel view synthesis.
@@
The work most closely related to ours is GPS-Gaussian [36], which proposes a stereo matching network for 3D Gaussian position estimation and employs two 3-layer U-Nets to predict 3D Gaussian scales, rotations, and opacities. Although GPS-Gaussian demonstrates the potential for real-time human reconstruction and novel view synthesis, it suffers from severe distortions under sparse camera settings and mismatch across multiple viewpoints. Subsequent works have attempted to alleviate these issues. For instance, Tele-Aloha [26] introduces an image blending and cascaded disparity estimation method for human reconstruction with four input views. However, this approach is tailored to a specific system and struggles to generalize to sparser camera settings. Although GHG [8] achieves real-time 3D Gaussian-based human novel view synthesis in a feed-forward manner, it requires additional human template priors, thus inheriting the limitations of template-based methods. In contrast, our method eliminates the need for human templates and is specifically designed to generalize effectively across diverse sparse camera settings.
@@ 3.1 Overview
where $\bm{\theta}$ denotes the learnable parameters of the network.
The framework of EVA-Gaussian is depicted in Fig. 2. EVA-Gaussian splits the process of predicting Gaussian maps into three stages. In the first stage, it employs a U-Net architecture with an Efficient cross-View Attention module (EVA) to obtain enhanced multi-view predictions of the 3D Gaussian position maps $\{\bm{P}_{i}\}_{i=1}^{n}$, as elaborated in Sec. 3.2. In the second stage, a Gaussian attribute prediction network, detailed in Sec. 3.3, takes the predicted position maps $\{\bm{P}_{i}\}_{i=1}^{n}$ and the original RGB images $\{\bm{I}_{i}\}_{i=1}^{n}$ as input to estimate the remaining attributes of 3D Gaussians. The predicted 3D Gaussians from all source images are then aggregated to render target views using differential rasterization [6]. In the final stage, the rendered image $\hat{\bm{I}}^{0}$ and its corresponding feature map $\bm{F}_{\text{novel}}$ are fused for further refinement using the network described in Sec. 3.4. In addition, an anchor loss is introduced during the training stage to enhance the overall reconstruction quality, as depicted in Sec. 3.5.
Figure 3: Efficient cross-View Attention (EVA) module for 3D Gaussian position estimation. EVA takes multi-view image features as input, embeds them into window patches using a shifted algorithm, and performs cross attention between the features from different views.
@@ 3.2 Gaussian Position Estimation
To tackle this challenge, we employ a U-Net based architecture, denoted by $\mathcal{D}^{P}_{\bm{\theta}_{1}}$, to estimate the 3D Gaussian position maps $\{\bm{P}_{i}\}_{i=1}^{n}$ from multi-view images $\{\bm{I}_{i}\}_{i=1}^{n}$. This process is expressed as:
@@
To ensure accurate depth estimation across diverse camera angles and arbitrary input views at high resolutions, while maintaining computational and temporal efficiency, we propose an EVA module, as illustrated in Fig. 3. This module is integrated into the three lowest resolution layers of the U-Net basebone $\mathcal{D}^{P}_{\bm{\theta}_{1}}$ to facilitate the multi-view correspondence retrieval and information exchange. We use $j$ to index these three layers, with $j=-1,j=-2,$ and $j=-3$ representing the lowest, second-lowest, third-lowest resolution layers, respectively. The EVA module takes intermediate image features $\bm{E}^{j}_{i}\in\mathbb{R}^{R^{j}\times C^{j}},\forall i\in\{1,\cdots,n\},\forall j\in\{-1,-2,-3\}$, as input, where $R^{j}$ and $C^{j}$ denote the total number of pixels and the channel dimension of each pixel at layer $j$, respectively. The module outputs enhanced image features $\widetilde{\bm{E}}^{j}_{i}$, which incorporate cross-view information. Before the execution of attention mechanisms, a learnable positional embedding $\gamma$ is added to encode spatial coordinates, improving the understanding of image geometry.
While cross-view attention mechanisms have been explored in various vision tasks that require multiple image inputs, existing approaches in feed-forward 3D Gaussian reconstruction, such as epipolar attention in pixelSplat [2] and vanilla cross-attention in MVSplat [4], are limited to low-resolution inputs (256$\times$256) and fail to recover fine-grained details. In contrast, our approach targets high-resolution (1024$\times$1024) human-centric reconstruction, a setting that has been demonstrated to be critical for recovering intricate human details [24]. The key novelty of our EVA module lies in its localized 1D window-based attention mechanism, which exploits the observation that corresponding pixels from reference views are typically located in adjacent positions along the x-axis in human-centric camera settings. Unlike global attention mechanisms used in prior works [4, 2, 12], EVA computes cross-attention only within a local 1D window aligned with the x-axis, and the window is shifted by half its size at each iteration to expand the receptive field. This design significantly reduces computational complexity and improves GPU memory utilization, while maintaining high performance, as demonstrated in Tab. 1.
Notably, when the scale of each Gaussian is sufficiently small, the 3D Gaussian position of a pixel aligns precisely with its corresponding value on the depth map. A detailed proof of this property is provided in Appendix D. Based on this observation, we train the position estimation network $\mathcal{D}^{P}_{\bm{\theta}}$ to obtain the position maps $\{\bm{P}_{i}\}_{i=1}^{n}$ with the mean squared error (MSE) loss function: